### v1

@@ Discussion
In this study, we introduced a KG-guided chain-of-thought framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data. By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we created a compact supervision corpus that supports data-efficient fine-tuning of lightweight LLMs.
Our results indicate that KG-guided CoT improves data-efficient prognostic prediction. On MIMIC-III visit-level disease prediction, KG-guided CoT fine-tuning substantially improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and brings them in line with strong classical baselines. With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$, matching or exceeding logistic regression and XGBoost while also producing explicit reasoning traces; with 1,000 visits, performance further improves to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$. These gains suggest that KG-anchored CoT supervision acts as an effective “reasoning prior,” helping small LLMs learn clinically meaningful decision boundaries from limited data. Classical models remain competitive on AUROC, but KG-guided LLMs tend to close or surpass this gap while improving macro-AUPR and F1, especially for minority conditions. Practically, this points toward hybrid designs where well-calibrated classical or deep risk models provide scores, and KG-guided LLMs serve as a clinician-facing layer that turns those scores into structured, patient-specific, KG-grounded explanations.
We also observe promising cross-cohort generalization and decision-boundary transfer. Our zero-shot transfer from MIMIC-III to CRADLE suggests that KG-guided CoT yields reasonably robust decision boundaries: without any retraining, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from about 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain relatively stable. This pattern indicates that overall discrimination is preserved, but the effective decision thresholds and class balance improve in the target cohort.
@@
Several limitations remain. Our approach depends on the coverage and quality of PrimeKG, so gaps or biases can propagate into CoT supervision. ICD–9-to-KG mapping and relevance/path selection rely on GPT-4o and heuristic thresholds, which may introduce subtle errors despite label-consistency filtering. Human annotators also highlighted opportunities to improve the input representation itself: our current summaries sometimes blur temporal structure, such as exposure versus outcome windows, and care setting (inpatient versus outpatient), which likely caps attainable performance and can lead even strong models to produce brittle or poorly justified reasoning. Finally, the current models require KG access at inference time for KG-conditioned variants.
Looking ahead, our results point to several concrete directions for future work. Applying the framework to richer EHR views, such as notes, medications, labs, and additional KGs could support more nuanced, multimodal reasoning, while replacing simple shortest-path guidance with richer graph reasoning may further improve explanation quality. Clinician feedback also points to a complementary direction: improving input structuring by making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and including basic demographics to reduce tendency for the model to hallucinate or over-interpret limited evidence. In parallel, future work may introduce temporal or guideline-based constraints to reduce illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
@@ Acknowledgements
This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University.
Appendix
Implementation Details
All fine-tuning experiments were conducted on a single NVIDIA H200 GPU on our institutional high-performance computing cluster, with 512 GB of host memory allocated via SLURM. We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT supervision corpus using full-parameter updates with AdamW (learning rate $1\times 10^{-5}$, cosine learning-rate schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096 tokens, and 10 training epochs in bf16 precision under DeepSpeed ZeRO-3. We monitored validation performance and selected the checkpoint with the best macro-AUPR. KG-guided CoT traces were generated using GPT-4o; across entity-mapping validation, node relevance selection, path pruning, and CoT generation, the total API cost for data creation was approximately \$100.
Prompt used:

### v2

@@ Discussion
In this study, we proposed a KG-guided chain-of-thought generation framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data. By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we construct a compact supervision corpus for data-efficient fine-tuning of lightweight LLMs. On MIMIC-III visit-level disease prediction, KG-guided CoT improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and makes them competitive with strong classical baselines. With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$; with 1,000 visits, performance increases to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$. These results suggest that KG-anchored CoT supervision acts as an effective reasoning prior, helping small LLMs learn clinically meaningful decision boundaries from limited data. Classical models remain strong on AUROC, but KG-guided LLMs close or surpass this gap while improving macro-AUPR. This points toward hybrid designs in which classical or deep risk models provide scores, and KG-guided LLMs act as a clinician-facing layer that produces explanations. We also observe encouraging cross-cohort generalization. In zero-shot transfer from MIMIC-III to CRADLE, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain stable. This pattern indicates that discrimination is preserved, while the induced decision boundaries and reasoning adapt to the target cohort.
@@
Several limitations remain. Our approach depends on the coverage and quality of PrimeKG, so gaps or biases can propagate into CoT supervision. The mapping of ICD–9 to KG and the selection of the relevance/path rely on GPT-4o and heuristic thresholds, which may introduce subtle errors during data generation.
Looking ahead, our results suggest several concrete directions. Extending the framework to richer EHR views (notes, medications, labs) and additional KGs could support more nuanced multimodal reasoning, while moving beyond simple shortest paths to richer graph reasoning may improve explanation quality. In addition, clinician feedback highlights the value of better input structuring—making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and adding basic demographics features (e.g. age, sex) —to reduce hallucinations and over-interpretation. Future work can introduce temporal or guideline-based constraints to further suppress illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
@@ Acknowledgements
This research was partially supported by internal funds and GPU servers provided by the Computer Science Department of Emory University.
Implementation Details.
All fine-tuning experiments were conducted on a single NVIDIA H200 GPU with 512 GB host memory on Emory’s high-performance computing cluster. We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT corpus using full-parameter AdamW (learning rate $1\times 10^{-5}$, cosine schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096, and 10 training epochs in bf16 under DeepSpeed ZeRO-3. We monitored validation performance and selected the checkpoint with the best macro-AUPR. KG-guided CoT traces were generated using GPT-4o; across entity-mapping validation, node relevance selection, path pruning, and CoT generation, the total API cost for data creation was approximately \$100. The exact prompt templates used in our pipeline are listed below.